[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Sep 29, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Fri Sep 29 16:32:10 UTC 2017


Here is today's summary of NHC computer operations:
--- Matt S.

----------

TAFB:
No new issues.
HSU:
1. NCEP model guidance (GFSI, HWFI, HMNI) were not available at 00z. Monica  
fixed a path in a script that was modified yesterday and then the models  
were available before the 03z advisories had to be finished.

2. Did not receive SHIPS/LGEM after submitting guidance for Maria, Lee, and  
AL99. Also, the SDM reported that they did not receive the inputs for the  
HWRF/HMON submissions. Monica investigated and determined that the compute  
file timestamp needed to be updated and the file resubmitted to fix both  
problems.

3. All N-AWIPS and ATCF workstations except sawfish and pac-atcf were  
rebooted yesterday.
Other items:
1. Doug Gaer started adding a number of NHC systems to NCO's Big Brother  
monitoring

NHC RFCs:
1. Monica Bozeman completed RFC 252...Upgrade ATCF FTP scripts to use rsync  
instead of scp.

2. Doug Gaer completed RFC 3684...NHC video card upgrades for HSU satellite  
display workstations.


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170929/afa54ec7/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list